soccer20:

Uri asked six friends how many different foreign countries they’d been to. He recorded 0, 1, 1, 2, 5, and 9 countries visited. What is the mean absolute deviation of countries visited? Round to the nearest tenth, if needed. 0 2.7 2.9 3

finny:

it can't be 0 and it can't be 3

soccer20:

ok

finny:

what do you think the answer is?

Narad:

The mean is \[\mu=(0+1+1+2+5+9)/6=18/6=3\] The mean absolute deviation is =( |0-3|+|1-3|+|1-3|+|2-3|+|5-3|+|9-3|)/6 =(3+2+2+1+2+6)/6 =16/6 =2.7